OpenStudy (anonymous):

Get a medal.................... Evaluate the expression when x = –18. x-10/4 A. 7 B. 2 C. –2 D. –7

OpenStudy (anonymous):

x = -18 \[\frac{ x-10 }{ 4 }\]

OpenStudy (opcode):

\[\dfrac{-18 - 10}{4}\] \[\dfrac{-28}{4}\] \[\dfrac{-28}{4}=-7\] So D. -7

OpenStudy (anonymous):

@Opcode Thanks, I thought it was D. too. Just wanted to make sure XD